# Phantom matter: a challenging solution to the cosmological tensions

### v1

@@ 5 Discussion and conclusions
All in all, model $w$XCDM offers not only a significantly better quality fit than model $\Lambda_{s}$, but at the same time does provide a richer conceptual frame for our understanding of the physical phenomena that are involved. As we have emphasized, the $w$XCDM is inspired in the old existing $\Lambda$XCDM framework[41, 42, 43], which is born in the theoretical arena of the RVM [8, 9]; and, at the same time, it combines very valuable elements of its low-energy string theory version[11, 10]. This is in stark contrast to the $\Lambda_{s}$ approach, where no obvious physical explanation is provided for the AdS-dS transition nor for its single occurrence precisely near our time. In the $w$XCDM, the formation of a PM bubble is induced by quantum fluctuations associated to the nearing of the universe towards a de Sitter phase. The phenomenon need not be unique: the bubbles of PM could well be operating more than once at earlier times on the same physical grounds, what would trigger an anomalous outgrowth of structures at even higher redshifts, say in the range $z\sim 5-10$. This might explain the appearance of the large scale structures recently spotted at unusually high redshifts by the JWST mission [30, 100, 101]. Such LSS ‘anomalies’, which find no explanation in the $\Lambda$CDM, could also be described within the current framework since they might be conceived as being the earliest bubbles of PM in the late universe, corresponding probably to the first tunneling attempts anticipating the eventual dS phase near our time. This is reasonable since the tunneling process towards the final de Sitter era should be gradual, exactly as in the early universe[11]. It is encouraging to realize that the physics involved in the PM scenario presented here for the solution of the cosmological tensions may ultimately stem from quantum phenomena, what gives real hope for eventually achieving a deeper connection of cosmology with fundamental physics.
@@
Acknowledgments This work is partially supported by grants PID2022-136224NB-C21 and PID2019-105614GB-C21, from MCIN/AEI/10.13039/501100011033. AGV is funded by “la Caixa” Foundation (ID 100010434) and the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 847648, with fellowship code LCF/BQ/PI21/11830027. JSP is funded also by 2021-SGR-249 (Generalitat de Catalunya) and CEX2019-000918-M (ICCUB, Barcelona). Both of us acknowledge networking support by the COST Association Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”.
Appendix: Breakdown of $\chi^{2}_{\rm min}$ contributions and full triangle plot

### v3

@@ 6 Conclusions
All in all, model $w$XCDM potentially offers a rich conceptual framework for our understanding of the physical phenomena that may be involved. In the $w$XCDM, the formation of PM bubbles is induced by quantum fluctuations associated to the nearing of the universe towards a de Sitter phase. The phenomenon need not be unique: the bubbles of PM could well be operating more than once at earlier times on the same physical grounds, what would trigger an anomalous outgrowth of structures at even higher redshifts, say in the range $z\sim 5-10$. This might explain the appearance of the large scale structures recently spotted at unusually high redshifts by the JWST mission (Labbé et al., 2023; Adil et al., 2023; Menci et al., 2024). Such LSS ‘anomalies’, which find no explanation in the $\Lambda$CDM, might also be described within the current proposal since they can be conceived as being the earliest bubbles of PM popping up in the late universe, corresponding probably to the first tunneling attempts anticipating the eventual dS phase near our time. This is reasonable, given the fact that the tunneling process towards the final de Sitter era should be gradual, exactly as in the early universe (Mavromatos & Solà Peracaula, 2021b).
Last but not least, as promised, we would like to advance here, if only very briefly, the outcome of the preliminary analysis performed with BAO 3D data. Our extended study has revealed that the $w$XCDM model with BAO 3D (replacing BAO 2D) data improves once more (and quite significantly) the overall fit to the cosmological data as compared to the $\Lambda$CDM, see (Gómez-Valent & Solà Peracaula, 2024). This should not come as a big surprise, if we take into account our discussion in the previous section, where we have shown by appealing to Table 2 in the Appendix that the composite DE models beat the $\Lambda$CDM in every single observable, not only in those driving the tensions. It means that the composite feature of the DE proves really fruitful to better accommodate the various sources of cosmological data. This is remarkable in the first place, for it means that the superiority of the composite DE models over the $\Lambda$CDM is not compromised at present by the particular type (2D or 3D) of BAO data used, despite them being currently under some tension (Camarena & Marra, 2020; Gómez-Valent et al., 2024; Favale et al., 2024b). Notwithstanding this relevant fact, which holds good for the global fit involving all sorts of cosmological data, we should not underemphasize that, insofar as concerns the $H_{0}$ tension itself, the fit with BAO 3D data proves unable to match the efficiency of the BAO 2D fit. Specifically, with BAO 3D the $H_{0}$ tension persists at a level of $\sim 2.9\sigma$ (no longer at the insignificant level of $0.25\sigma$ like in the BAO 2D case). Thus, while it is true that with BAO 3D the $H_{0}$ tension gets somewhat reduced from the large initial value ($\sim 5\sigma$ c.l.), it remains still sizeable. In contrast, we find that the growth tension is less affected. All in all, a final solution to the cosmological tensions – above all to the $H_{0}$ tension – is still being compromised by a lack of consistency between the two sources of BAO, angular and anisotropic. The systematic exposition of the results with BAO 3D, along with the quantitative differences with respect to the analysis using BAO 2D, is beyond the scope of this work and its presentation is left for a separate publication (Gómez-Valent & Solà Peracaula, 2024).
At the end of the day, it is encouraging to realize that the physics involved in the PM scenario presented here for a potential solution to the cosmological tensions, might ultimately stem from (dynamical) quantum vacuum phenomena, what gives hope for eventually achieving a deeper understanding of the cosmological evolution of the universe from fundamental physics. Whether this is actually the case or not will depend, of course, on the eventual resolution of the BAO tension itself, which appears to be the main stumbling block for the ultimate understanding of the problem.
7 Acknowledgements
@@
This work is partially supported by grants PID2022-136224NB-C21 and PID2019-105614GB-C21, from MCIN/AEI/10.13039/501100011033. AGV is funded by “la Caixa” Foundation (ID 100010434) and the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 847648, with fellowship code LCF/BQ/PI21/11830027. JSP is funded also by 2021-SGR-249 (Generalitat de Catalunya) and CEX2019-000918-M (ICCUB, Barcelona). Both of us acknowledge networking support by the COST Association Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”.
Full triangle plot and breakdown of $\chi^{2}_{\rm min}$ contributions